OpenStudy (anonymous):

Can you help me figure this out @mahmit2012,@heena,@eliassaab, @arvaindG.

OpenStudy (anonymous):

Explain the law of motion needed to solve this problem and then solve it. On Planet Zorg, a 30 kg barbell can be lifted by only exerting a force of 180 N. What is the acceleration of gravity on Planet Zorg? Predict how the motion of this object would change if on Earth.

OpenStudy (anonymous):

@mahmit2012 @heena @eliassaa @arvaindG

OpenStudy (btaylor):

From Newton's second law (F=ma): a = F/m = 189 N/30 kg = 6.0 m/s2

OpenStudy (anonymous):

acceleration=force/mass use it u ll get acceleration on the planet

OpenStudy (anonymous):

thanks!

OpenStudy (btaylor):

thank google.

OpenStudy (anonymous):

lol